Cherry-red embers in a fireplace are at 875∘C and have an exposed area of 0.2 m2 and an emissivity of 1.0. The surrounding room has a temperature of 21∘C. If 50% of the radiant energy enters the room, what is the net rate of radiant heat transfer, into the room, in kilowatts?
Answer #1


given
T1 = 875 C = 875 + 273 = 1148 K
T2 = 21 C = 21 + 273 = 294 K
A = 0.2 m^2

using Stefan-Boltzman's law,
the rate of radiant energy enters the room = 0.5*A*sigma*(T1^4 - T2^4) (here sigma is stefan's constant)

= 0.5*0.2*5.67*10^-8*(1148^4 - 294^4)

= 9806 W

= 9.806 kW <<<<<<<<---------------Answer